OpenStudy (anonymous):

Precal HELP please, Rational equations graphing list all intercepts and all asymptotes using correct format 1) f(x) = 1/x 2) g(x) = 4x +3 3 / x+4

jimthompson5910 (jim_thompson5910):

To find the y intercept, plug in x = 0 and evaluate f(x) = 1/x f(0) = 1/0 ... replace every x with 0 f(0) = undefined it's undefined because we cannot divide by zero. So there are NO y intercepts

jimthompson5910 (jim_thompson5910):

To find the x intercepts, you replace f(x) with 0 and solve for x f(x) = 1/x 0 = 1/x ... replace f(x) with 0 what do you get when you solve for x?

OpenStudy (anonymous):

0? and thank you so much!